Three months ago, Jane had five times as many DVDs as Duncan. Since then they have bought 12 more DVDs Both Jane now has twice as many as Duncan. How many DVDs does Jane have now?
Thank you

Answer by rothauserc(4718)

Let Duncan had x DVDs 3 months ago and Jane had 5x
they each bought 12 and now Jane has twice as many as Duncan
5x+12 = 2*(x+12) and
5x+12 = 2x+24 gives
3x = 12 and x =4
Jane now has 2*(4+12) or 32 DvDs